# Securing Automated Insulin Delivery Systems: A Review of Security Threats and Protectives Strategies

### v1

@@ Abstract
Automated insulin delivery (AID) systems have emerged as a significant technological advancement in diabetes care. These systems integrate a continuous glucose monitor, an insulin pump, and control algorithms to automate insulin delivery, reducing the burden of self-management and offering enhanced glucose control. However, the increasing reliance on wireless connectivity and software control has exposed AID systems to critical security risks that could result in life-threatening treatment errors. This review first presents a comprehensive examination of the security landscape, covering technical vulnerabilities, legal frameworks, and commercial product considerations, and an analysis of existing research on attack vectors, defence mechanisms, as well as evaluation methods and resources for AID systems. Despite recent advancements, several open challenges remain in achieving secure AID systems, particularly in standardising security evaluation frameworks and developing comprehensive, lightweight, and adaptive defence strategies. As one of the most widely adopted and extensively studied physiologic closed-loop control systems, this review serves as a valuable reference for understanding security challenges and solutions applicable to analogous medical systems.
keywords:
automated insulin delivery systems, artificial pancreas systems , security , cybersecurity , attack vectors, defence , physiologic closed-loop control system
@@
[label1]organization=College of Computing and Data Science, Nanyang Technological University,addressline=50 Nanyang Ave, postcode=639798, country=Singapore
1 Introduction
The management of Type 1 diabetes (T1D), a chronic condition where pancreatic beta cells produce insufficient insulin for regulating blood glucose (BG) levels, remains a significant challenge for millions of individuals worldwide [1]. As there is no realistic biological cure in the near future, the treatment primarily involves intensive manual insulin therapy to manage BG levels in the range to prevent hyperglycemia or hypoglycemia, both of which can lead to severe, life-threatening complications [2]. However, this treatment regimen requires frequent BG monitoring and insulin adjustment, which are time-consuming and error-prone. To address these challenges, automated insulin delivery (AID) systems, also known as artificial pancreas (AP) systems, have emerged as a significant technological advancement in diabetes care [3].
AID system is a wearable physiological closed-loop control system that integrates a continuous glucose monitor (CGM), an insulin pump, and control algorithms to modulate insulin delivery based on continuous glucose measurements. This system not only alleviates the burden of self-management but also significantly improves clinical outcomes [4]. The demand for AID systems surged by advanced technology and the growing global prevalence of diabetes [5][6]. In 2023, the global AID systems market reached US\$ 2,338.5 million and is projected to expand to US\$ 5,800.2 million by 2032 [7]. However, the complex interconnectivity, increasing reliance on automated control, and wireless connectivity for wearable convenience in AID systems introduce significant risks that can lead to real-time treatment errors with potentially life-threatening consequences.
Threats to the reliability and safety of AID systems primarily arise from two sources: unintended failures (i.e., device limitations and malfunctions) and malicious attacks. Unintended failures stem from device physiological, technological, and behavioural limitations [3], as well as system malfunctions [8]. For instance, physiological constraints on delayed insulin absorption and postprandial rapid glucose fluctuations can render ineffective insulin delivery of AID systems. Technological limitations, including inaccurate glucose readings due to noise spikes, signal dropouts, miscalibration, or pressure applied to the sensor site (e.g., lying on the sensor while sleeping), can result in inappropriate insulin dosing. Additionally, behavioural limitations, such as improper patient operation, and malfunctions in critical components like infusion site failures and cartridge errors, can also lead to AID systems failures.
Malicious attacks, on the other hand, involve adversaries exploiting vulnerabilities in AID systems to intercept patient data or impact insulin delivery. These cybersecurity threats have escalated with the adoption of wireless communication in AID systems [9]. Examples include unauthorised remote control [10] and insulin dose manipulation attacks [11]. Early experiments have shown the feasibility of cyberattacks through unencrypted wireless channels to deliver lethal insulin doses from at least 20 meters away [12][13]. The U.S. Food & Drug Association (FDA) issued warnings in 2019 regarding AID systems cybersecurity weaknesses [14], and commercial AID systems products have been recalled due to security concerns [15].
@@ 1 Introduction
Review Paper Focus Discussion on Unintended Failures Discussion on Malicious Attacks Analysis of Attack Vectors Evaluation of Defence Mechanism Klonoff, 2015 [16] Cybersecurity for Connected Diabetes Devices $\times$ $\checkmark$ $\times$ (no literature review) O’Keeffe et al,. 2015 [17] Cybersecurity in AID systems $\times$ $\checkmark$ $\times$ (no literature review) Blauw et al., 2016 [18] Safety and design requirements of the AID systems $\checkmark$ $\times$ $\times$ (For unintended failures) Ramkissoon et al., 2017 [19] Specific hazards applicable to the AID systems $\checkmark$ $\times$ $\times$ (For unintended failures) Thompson et al., 2018 [20] Safety issues in outpatient settings $\checkmark$ $\checkmark$ $\times$ $\times$ Quintal er al., 2019 [21] Ethical issues associated with AID systems $\checkmark$ $\checkmark$ $\times$ $\times$ Singh and Rahman, 2020 [22] The security of Open AID systems system $\times$ $\checkmark$ (no literature review) (no literature review) Sherr et al., 2022 [3] Current landscape and safety challenges of AID systems $\checkmark$ $\checkmark$ $\times$ $\times$ Wirth, 2022 [23] The states and tendency of cybersecurity in the diabetes ecosystem $\times$ $\checkmark$ $\times$ $\times$ Nazzal et al., 2024 [24] Software safety and security verification and validation on AID systems $\times$ $\times$ $\times$ (Formal Verification methods) Kapadia, 2024 [25] Limitations and malfunctions $\checkmark$ $\times$ $\times$ $\times$ This paper The cybersecurity status, attack vectors, and proposed defence strategies of AID systems $\times$ $\checkmark$ $\checkmark$ $\checkmark$
Table 1: Comparison of review papers on the security of AID systems in Last 10 Years.
@@
Strengthening AID systems security requires a thorough understanding of existing and emerging threats, a critical evaluation of defence mechanisms, and the identification of gaps that must be addressed. As summarized in Table 1, most review papers published in the past decade have primarily focused on unintended failures [18] [19] [25] or discussed threats from malicious attacks without analysing specific attack vectors and defence mechanisms [26] [21] [3]. While several studies [16] [17] [27] [28] provided general overviews of cyber vulnerabilities and potential countermeasures, they lack in-depth examinations of related studies. Nazzal et al. [25] conducted a thorough review of software verification methods for AID systems; however, broader security defence strategies remain unexplored.
@@
Besides, compared to most other safety-critical devices within the Internet of Medical Things (IoMT), which also face threats from malicious attacks, AID systems present unique security challenges: (1) the human-in-the-loop component, which introduces individual physiological variability and unpredictable external factors, especially in free-living settings, and (2) constrained power, computational, and memory resources, which limit the feasibility of typical resource-intensive security strategies. These challenges not only amplify the impact of security threats but also complicate the development of effective defence mechanisms. While existing IoMT security reviews [29][30][31] discussed AID systems cybersecurity, they fail to address its unique risks. Notably, no prior reviews comprehensively examine security threats and defence strategies for wearable physiologic closed-loop control systems.
@@
To address this gap, this study systematically examines the security of AID systems, providing a detailed analysis of attack vectors, existing defence strategies, and potential areas for future research. We introduce a taxonomy that categorizes security threats and defence mechanisms in AID systems, as shown in Figure 2. Given the increasing adoption of AID systems and the growing body of research in this field, this review offers valuable insights for enhancing the security of similar medical systems.
@@
The paper is organized as follows (see Figure 1). Section 3 explores AID systems cybersecurity from technical, legal, and commercial perspectives, establishing a foundational understanding of the field. Section 4 examines existing and emerging attack vectors, followed by Section 5, which reviews existing defence strategies. Section 6 discusses current evaluation methods and available research resources. Finally, Section 7 summarizes key findings and identifies research gaps, and Section 8 outlines future directions for strengthening AID systems security.
@@
2 Methodology and Materials
To comprehensively assess the security landscape of AID systems, this review adopts a hybrid methodology, integrating elements of both survey and systematic review. The review begins with a broad investigation into AID systems cybersecurity, covering technical vulnerabilities, regulatory frameworks, and security measures in commercial products. The materials considered extend beyond peer-reviewed journals and conference papers to include diverse sources such as clinical studies, industry reports, regulatory standards, product manuals, and relevant online resources. Following this, the PRISMA framework [32] was employed for systematic analysis of attack vectors and defence mechanisms for AID systems security. To guide this analysis, the following research questions (RQs) were formulated:
  • 1.
@@ 2 Methodology and Materials
A systematic literature search was conducted across multiple databases, including Scopus, Google Scholar, Web of Science, IEEE Xplore, and PubMed. Boolean operators and keyword variations were applied, as summarized in Table 2. The review focused on publications from 2010 to the collection date (February 11, 2025), with studies outside this range filtered out during the initial search.
@@
  • 3.
    Studies proposing attack vectors, defence mechanisms, or risk evaluations for artificial pancreases
@@
From the initial database search, 53 relevant papers were identified. To ensure comprehensive coverage, backward and forward citation tracking was conducted, yielding 23 additional studies. These were initially missed due to: (1) Broad scope—studies focusing on medical devices, IoT-based healthcare systems, or cyber-physical systems that included AID systems as a case study; and (2) Terminology differences—studies referring to AID systems under alternative terms such as wireless insulin pump systems or sensor-augmented pumps.
@@
3 Background
3.1 Technical Description and Vulnerabilities
@@
Figure 5: An overview of a hybrid closed-loop insulin delivery system, illustrating core components, workflow and potential functional vulnerabilities. Dotted lines indicate wireless communication links between modules, while solid lines represent direct attachments to the patient’s body or integrated elements - adapted from [33].
@@
AID combines a CGM, an insulin pump and a controller algorithm to calculate and deliver suitable insulin doses, helping to maintain blood glucose levels within the target range. Figure 5 shows the workflow of a hybrid closed-loop insulin delivery system, the paradigm common among current FDA-approved AID systems products. In this system, the CGM periodically (e.g., every 5 minutes) transmits the measured glucose levels to the control algorithm device, which calculates basal rates and corrective bolus doses of insulin, and then directs the subcutaneous insulin pump to deliver the specified dose into the patient body [34]. Unlike fully closed-loop systems that automate all insulin delivery, hybrid systems still require user input for meal and activity-related boluses [35]. Fully closed-loop systems, though promising, still face challenges in managing postprandial hyperglycemia and exercise-related glycemic fluctuations [36]. Despite different automation levels, both hybrid and fully closed-loop systems share core modules and rely on wireless connectivity and software control, exposing them to similar cybersecurity vulnerabilities.
@@
In the context of AID systems, cybersecurity focuses on protecting digital assets, which encompass physiological data, treatment commands transmitted wirelessly between system components, the software controlling these processes, and the patient’s personal information stored within the device. Vulnerabilities exist both in wireless communication channels and individual system modules, as detailed in the following subsections.
3.1.1 Wireless communication
@@
The communication between the three main components of AID systems primarily relies on wireless transmissions for wearable convenience. Safe information exchange requires secure, lightweight, and resilient communicational protocols, including proper authentication and encryption. However, as examined in [37], applied protocols often fall short of meeting safety requirements. For example, the widely used Bluetooth LE protocol is vulnerable to replay and man-in-the-middle attacks [38], allowing adversaries to eavesdrop on communications or inject faulty glucose data and erroneous control commands.
3.1.2 CGM
@@
A CGM system continuously tracks blood glucose (BG) levels in the patient’s body and transmits the collected physiological data to the controller every few minutes. It consists of a disposable glucose sensor, placed under the skin to measure glucose levels, and a transmitter attached to the sensor that sends the data to the control algorithm platform. Accurate BG readings are crucial for safe insulin delivery, as any incorrect values can potentially result in harmful dosing decisions. However, a compromised CGM system may operate maliciously by providing inaccurate BG levels, either due to device tampering or data manipulation during wireless transmission [12][34][13].
3.1.3 Control Algorithms
@@
The controller computes the recommended insulin doses and transmits infusion commands to the insulin pump. Its decision-making process is based on real-time sensor data, historical records, and estimated patient status (e.g., eventual BG, Insulin on Board (IOB)). The controller has a patient physiological model to estimate glucose-insulin dynamics in the patient’s body, and control algorithms to determine optimal insulin doses. These algorithms are typically deployed on smartphones, tablets, personal computers, or integrated with insulin pumps, which are often the platforms for user monitoring of device activities. Three main security threats are associated with the controller:
  • 1.
@@
3.1.4 Subcutaneous Insulin Pump

### v2

# Securing Automated Insulin Delivery Systems: A Review of Security Threats and Protective Strategies

@@ Abstract
Automated Insulin Delivery (AID) systems represent a significant advancement in diabetes care and wearable physiological closed-loop control technologies, integrating continuous glucose monitoring, control algorithms, and insulin pumps to improve blood glucose level control and reduce the burden of patient self-management. However, their increasing dependence on wireless communication and automatic control introduces security risks that may compromise patient privacy or result in life-threatening treatment errors. This paper presents a comprehensive survey of the AID system security landscape, covering technical vulnerabilities, regulatory frameworks, and commercial security measures. In addition, we conduct a systematic review of attack vectors and defence mechanisms proposed in the literature, following the PRISMA framework. Our findings highlight critical gaps, including the lack of specific security evaluation frameworks, insufficient protections in real-world deployments, and the need for comprehensive, lightweight, and adaptive defence mechanisms. We further investigate available research resources and outline open research challenges and future directions to guide the development of more secure and reliable AID systems. By focusing on AID systems, this review offers a representative case study for examining and improving the cybersecurity of safety-critical medical wearable systems.
I Introduction
The management of Type 1 Diabetes (T1D), a chronic condition where pancreatic beta cells produce insufficient insulin for regulating blood glucose (BG) levels, remains a significant challenge for millions of individuals worldwide [37]. As there is no realistic biological cure in the near future, the treatment primarily involves intensive insulin therapy to maintain BG levels within a safe range, thereby mitigating the risk of serious complications caused by hyperglycemia or hypoglycemia [51]. However, this regimen of frequent BG measurement and insulin adjustments is time-consuming and error-prone for patient self-management. To address these challenges, automated insulin delivery (AID) systems, also known as artificial pancreas, have emerged as a promising “technical” solution [118].
AID system integrates a continuous glucose monitor (CGM), a control algorithm, and an insulin pump to automatically adjust insulin delivery based on real-time glucose measurements. With recent technological advancements, these systems have evolved into networked and intelligent medical wearables that improve glycemic outcomes and quality of life for individuals with diabetes [26][40]. Nevertheless, the increasing device complexity, coupled with their reliance on wireless connectivity and automated control, introduces significant safety and security concerns [6]. Ensuring the trustworthiness and resilience of AID systems against a broad spectrum of threats is essential for safeguarding patient health and enabling wider clinical adoption.
Threats to the reliability of AID systems primarily arise from two sources: unintended failures and malicious attacks. Unintended failures result from system limitations [118] and malfunctions [67]. For example, physiological constraints on delayed insulin absorption and postprandial rapid glucose fluctuations can limit the effectiveness of insulin delivery. Technological and behavioural limitations, such as sensor noise, signal dropouts, miscalibration, or pressure-induced sensor attenuation (e.g., lying on the sensor during sleep), may lead to inaccurate glucose readings and inappropriate dosing. Additionally, malfunctions with CGM and insulin pumps (e.g., early sensor failure, occlusion of insulin infusion sets, and cartridge errors) can further compromise system performance.
Malicious attacks, on the other hand, involve adversaries exploiting vulnerabilities in AID systems to intercept sensitive patient data or manipulate insulin delivery. These risks have escalated with the adoption of wireless communication in AID systems [61]. Demonstrated attacks include unauthorised remote control [84] and false CGM reading injection [112]. Early experiments have shown the feasibility of intercepting communication and launching lethal insulin doses through unencrypted wireless channels at least 20 meters away [73][27]. Recognising these risks, the U.S. Food & Drug Administration (FDA) issued warnings in 2019 regarding cybersecurity weaknesses in AID systems [3], and several commercial products have been recalled due to security concerns [115].
TABLE I: Comparison of review papers addressing AID system security over the past decade.
@@ I Introduction
Review Paper Focus Discussion on Unintended Failures Discussion on Malicious Attacks Analysis of Attack Vectors Analysis of Protective Mechanisms Klonoff, 2015 [66] Cybersecurity for Connected Diabetes Devices $\times$ $\checkmark$ $\times$ (No literature review) O’Keeffe et al,. 2015 [98] Cybersecurity in AID systems $\times$ $\checkmark$ $\times$ (No literature review) Blauw et al., 2016 [18] Safety and design requirements of the AID systems $\checkmark$ $\times$ $\times$ (Only for unintended failures) Ramkissoon et al., 2017 [111] Specific hazards applicable to the AID systems $\checkmark$ $\times$ $\times$ (Only for unintended failures) Thompson et al., 2018 [125] Safety issues in outpatient settings $\checkmark$ $\checkmark$ $\times$ $\times$ Quintal er al., 2019 [109] Ethical issues associated with AID systems $\checkmark$ $\checkmark$ $\times$ $\times$ Singh and Rahman, 2020 [119] The security of Open AID systems system $\times$ $\checkmark$ (No literature review) (No literature review) Sherr et al., 2022 [118] Current landscape and safety challenges of AID systems $\checkmark$ $\checkmark$ $\times$ $\times$ Wirth, 2022 [141] The states and tendency of cybersecurity in the diabetes ecosystem $\times$ $\checkmark$ $\times$ $\times$ Nazzal et al., 2024 [5] Software safety and security verification and validation on AID systems $\times$ $\times$ $\times$ (Only formal verification methods) Kapadia, 2024 [59] Limitations and malfunctions $\checkmark$ $\times$ $\times$ $\times$ This paper The cybersecurity status, attack vectors, and proposed defence strategies of AID systems $\times$ $\checkmark$ $\checkmark$ $\checkmark$
@@
Strengthening the reliability of AID systems requires a thorough understanding of existing and emerging threats, a critical evaluation of defence mechanisms, and the identification of gaps that must be addressed. Review papers on the reliability of AID systems published in the past decade, as summarized in Table I, most primarily focused on discussing unintended failures [18] [111] [59] or discussing threats from malicious attacks without analysing specific attack vectors and defence mechanisms [118][125] [109]. While several studies provided general overviews of cyber vulnerabilities and potential countermeasures, they lack in-depth examinations of related studies [66] [98] [119][141]. Nazzal et al. [5] conducted a thorough literature review of software verification methods for AID systems; however, broader security defence strategies remain unexplored. Besides, while some existing reviews on Internet of Medical Things (IoMT) security briefly discuss AID systems, they do not adequately address their distinctive risks [139][45][57].
@@
Compared to most other safety-critical IoMT systems, AID systems present unique security challenges. First, the physiological closed-loop control introduces substantial interpatient variability and unpredictable external factors, complicating threat detection and leaving a narrow window for preventing attacks from impacting drug delivery. Second, AID wearables operate under stringent power, computation, and memory constrained, limiting the feasibility of resource-intensive security strategies. These challenges not only amplify the impact of potential threats but also complicate the development of effective defence mechanisms. Notably, no prior reviews comprehensively examine security threats and protective strategies for wearable physiologic closed-loop control systems.
@@
To address this gap, this study systematically reviews the security of AID systems, providing an in-depth analysis of attack vectors, existing defence strategies, and open research challenges. We introduce a taxonomy that categorises security threats and defence mechanisms in AID systems, as illustrated in Figure 2. With the growing adoption of AID technologies and the expanding body of related research, this review not only advances understanding of AID security but also serves as a representative case study for strengthening the cybersecurity of other medical wearable systems.
@@
The remaining paper is organised as follows (see Figure 1). Section II presents the methodology and materials used in the review. Section III examines the security status of AID systems from technical, legal, and commercial perspectives, establishing a foundational understanding of the field. Section IV reviews existing and emerging attack vectors, and Section V analyses defence strategies proposed in the literature. Section VI discusses current evaluation methods and available research resources. Section VII discusses key findings and research gaps, and Section VIII outlines future directions for strengthening the security of AID systems.
@@
II Methodology and Materials
To comprehensively assess the security landscape of AID systems, this paper adopts a hybrid methodology, integrating elements of both survey and systematic review. The review begins with a broad investigation into AID systems cybersecurity, covering technical vulnerabilities, regulatory frameworks, and security measures in commercial products. The materials considered extend beyond peer-reviewed journals and conference papers to include diverse sources such as clinical studies, industry reports, regulatory standards, product manuals, and relevant online resources. Following this, the PRISMA framework [100] was employed for a systematic review of attack vectors and defence mechanisms proposed in the literature for AID systems security. To guide this analysis, the following research questions (RQs) were formulated:
@@ II Methodology and Materials
A systematic literature search was conducted across multiple databases, including Scopus, Google Scholar, Web of Science, IEEE Xplore, and PubMed. Boolean operators and keyword variations were applied, as summarized in Table II. The review focused on publications from 2010 to the collection date (February 11, 2025), with studies outside this range filtered out during the initial search.
TABLE II: Keywords used in search engines.
@@
  • Studies proposing attack vectors, defence mechanisms, or risk evaluations for AID systems
@@
From the initial database search, 53 relevant papers were identified. To ensure comprehensive coverage, backward and forward citation tracking was conducted, yielding 23 additional studies. These were initially missed due to: (1) Broad scope—studies focusing on medical devices, IoT-based healthcare systems, or cyber-physical systems that take AID systems as a case study; and (2) Terminology differences—studies referring to AID systems under alternative terms such as wireless insulin pump systems or sensor-augmented pumps.
@@
III Background
III-A Technical Description and Vulnerabilities
@@
Figure 5: An overview of a hybrid closed-loop insulin delivery system, illustrating core components, workflow, and potential functional vulnerabilities. Dotted lines indicate wireless communication links between modules, while solid lines represent direct attachments to the patient’s body or integrated elements - adapted from [77].
@@
AID combines a CGM, an insulin pump, and a controller algorithm to calculate and deliver suitable insulin doses, helping to maintain blood glucose levels within the target range. Figure 5 shows the workflow of a hybrid AID system, the paradigm common among current available products. In this system, the CGM periodically (e.g., every 5 minutes) transmits the measured glucose levels to the control algorithm device, which calculates basal insulin and corrective boluses in response to sensor glucose values and meal registration, and then directs the subcutaneous insulin pump to deliver the specified dose into the body [13]. Unlike full AID systems that automate all insulin delivery, hybrid systems require the user to administer prandial insulin before or at the time of meals to control post-meal blood glucose spikes [48]. Full AID systems, though promising, still face challenges in managing postprandial hyperglycemia and exercise-related glycemic fluctuations [124]. Despite different automation levels, both hybrid and full AID systems share core modules and rely on wireless connectivity and software control, exposing them to similar cybersecurity vulnerabilities.
@@
Cybersecurity concerns the protection of digital assets. In the context of AID systems, these assets include physiological and treatment data transmitted wirelessly between system components, the software controlling processes, and patient information stored on the device. Vulnerabilities may arise in both wireless communication channels and individual system modules, as detailed in the following subsections.
III-A1 Wireless communication
@@
Communication among the three core components of AID systems primarily relies on wireless transmissions for wearable convenience, making it possible for attackers to intercept, disrupt, or manipulate signals from a distance. Secure information exchange requires communication protocols that provide robust encryption, authentication, and resilience under emergencies. However, as examined in [64], protocols commonly applied in AID systems are often lightweight but fail to meet these security requirements. For example, the widely used Bluetooth Low Energy (BLE) protocol is vulnerable to replay and man-in-the-middle attacks, allowing adversaries to eavesdrop or inject faulty glucose readings and control commands [34].
III-A2 CGM
@@
A CGM system continuously tracks BG levels in the patient’s body and transmits the collected physiological data to the controller every few minutes. It consists of a disposable glucose sensor, placed under the skin to measure glucose levels, and a transmitter attached to the sensor that sends the data to the control algorithm platform. Accurate BG readings are crucial for safe insulin delivery, as any incorrect values can potentially result in harmful dosing decisions. However, a compromised CGM system may operate maliciously by providing inaccurate BG levels, either due to device tampering or data manipulation during wireless transmission [73][27][13].
III-A3 Control Algorithms
@@
The controller computes the recommended insulin doses and transmits infusion commands to the insulin pump. It incorporates patient physiological models to estimate glucose-insulin dynamics and control algorithms to determine optimal insulin doses. These computations rely on real-time sensor data, historical records, and estimated patient status (e.g., eventual BG, Insulin on Board (IOB)). Control algorithms may be deployed on smartphones, tablets, personal computers, or integrated into insulin pumps, which are often the platforms for user monitoring device activities. Three main security threats are associated with the controller:
@@
III-A4 Subcutaneous Insulin Pump